OpenStudy (anonymous):

A line with a slope of -1/20 passes thrpugh the point (20,19). What is its equation in slope-intercept form?

OpenStudy (anonymous):

y = m x + b y=-1/20, x=20 and y=19 from the slope and point coordinates respectively. solve for b. The line equation is: \[y=-\frac{1}{20} x+20\]

OpenStudy (anonymous):

Sorry. My bad. m = -1/20, not y = -1/20